# Maxitive functions with respect to general orders

### v2

@@ 4.5. Dispersive stochastic order
Remark 4.12.
The dispersive order is not antisymmetric and is therefore not a partial order. However, by considering the quotient space $\mathcal{M}/_{\sim}$ w.r.t. the equivalence relation $\xi{\sim}\eta$ if and only if $\xi=\eta+c$ for some $c\in\mathbb{R}$, we obtain that the partially ordered space $(\mathcal{M}/_{\sim},\preccurlyeq_{\rm disp})$ is a lattice, see [17]. It is shown in Corollary C.4 that the lattice $(\mathcal{M}_{\infty}/_{\sim},\preccurlyeq_{\rm disp})$ is order complete.
It follows from Proposition C.3 that every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{\infty}$ which is bounded from above w.r.t. $\preccurlyeq_{\rm disp}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determined by
@@
In the following, we consider the family $\Phi_{\rm disp}:=\{\phi_{u,v}\colon 0<u<v<1\}$ of functions $\phi_{u,v}\colon\mathcal{M}_{\infty}\to\mathbb{R}$ defined by $\phi_{u,v}(\xi):=q_{\xi}(v)-q_{\xi}^{+}(u)$. We identify $\Phi_{\rm disp}$ with the set $\Delta:=\{(u,v)\in\mathbb{R}^{2}\colon 0<u<v<1\}$ which we endow with the Euclidean topology. [^7] Although $\Delta$ is not a complete metric space, it is completely metrizable as it is an open subset, see [27, Theorem 24.12]. In particular, $\Delta$ is a Polish space with the Euclidean topology. It follows from (4.1) that the family $\Phi_{\rm disp}$ induces the dispersive order on $\mathcal{M}_{\infty}$. By Proposition B.1, the mapping $(u,v)\mapsto\phi_{u,v}(\xi)$ is lower semicontinuous on $\Delta$ for all $\xi\in\mathcal{M}_{\infty}$. Hence, the ordered space $(\mathcal{M}_{\infty},\preccurlyeq_{\rm disp})$ satisfies the assumptions of Section 3.
@@
By Proposition C.3, the sequence $(\xi_{n})_{n\in\mathbb{N}}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determind by
@@
Let $I:=\{s\in\mathbb{R}\colon A_{s}\neq\emptyset\}$. If $\psi$ has countably directed sublevel sets, it follows from Lemma 4.14 that for every $s\in I$, the set $Q_{s}:=\{q_{\xi}\colon\xi\in A_{s}\}$ is of bounded variation, and $\alpha_{\min}(s,u,v)=\beta(s,v)-\beta(s,u)$ for all $(u,v)\in\Delta$ where $\beta(s,u):={\rm TV}_{[0,u]}(Q_{s})$. Moreover, the mapping $\beta\colon\mathbb{R}\times(0,1)\to\mathbb{R}$ is non-decreasing in the second argument and satisfies $\beta(s,v)-\beta(s,u)\leq\beta(s^{\prime},v)-\beta(s^{\prime},u)$ for all $s\leq s^{\prime}$. As a direct application of Theorem 3.1, we obtain the following representation result for maxitive functions w.r.t. the dispersive stochastic order.
Theorem 4.15.
Let $\psi\colon\mathcal{M}_{\infty}\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$. Then the following conditions are equivalent:
@@
    $\psi$ is monotone and $(A_{s})_{s\in\mathbb{R}}$ are countably directed upwards w.r.t. $\preccurlyeq_{\rm disp}$.
@@ Appendix A Auxiliary results for maxitive risk measures on $C_{b}(S)$
The function $\psi$ is continuous from below if $\psi(f)=\lim_{n\to\infty}\psi(f_{n})$ for all sequences $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{X}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ and $f=\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{X}$. Moreover, the function $\psi$ is countably maxitive if $\psi(f)=\sup_{n\in\mathbb{N}}\psi(f_{n})$ for all sequences $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{X}$ with $f=\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{X}$.
@@
$\psi^{\ast}_{C_{b}(S)}(\mu):=\sup_{f\in C_{b}(S)}\left\{\int f\,{\rm d}\mu-\psi(f)\right\}.$
Here, ${\rm ca}_{1}^{+}(S)$ denotes the set of all Borel probability measures on $S$. In the particular case of a Dirac measure $\delta_{x}$, we get
$I_{\min}(x):=\psi^{\ast}_{C_{b}(S)}(\delta_{x})=\sup_{f\in C_{b}(S)}\{f(x)-\psi(f)\quad\mbox{for all }x\in S.$
The following result was recently shown by Delbaen [4, 5], see also [15] for an alternative proof.
@@
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T), (M) and (C). Then the following statements are equivalent:
@@
In combination with [12, Theorem 2.1], we obtain the following representation result.
@@
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T) and (M). Then the following statements are equivalent:
@@
    $\psi$ is countably maxitive.
@@
    $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$ for every sequence $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$.
@@
The argumentation is analogous to that in [12, Theorem 2.1]. Similar to the proof of [12, Theorem 2.1], it holds
@@
for all $\mu\in{\rm ca}_{1}^{+}(S)$, where ${\rm ess.sup}_{\mu}\mathcal{A}$ denotes the essential supremum of $\mathcal{A}$ w.r.t. $\mu$.
$(i)\Rightarrow(ii)$: Let $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since $\psi$ is countably maxitive, it holds $\psi(\sup_{n\in\mathbb{N}}f_{n})=\sup_{n\in\mathbb{N}}\psi(f_{n})\leq 0$, which shows that $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$.
$(ii)\Rightarrow(i)$: Let $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since by (T), $f_{n}-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$ for all $n\in\mathbb{N}$, we obtain $f-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$. By (T) and (M), it holds
@@
$(i)\Rightarrow(iv):$ Since $\psi$ is countably maxitive, by [12, Proposition 2.1], it satisfies (C). Fix $\mu\in{\rm ca}_{1}^{+}(S)$. Since $\mathcal{A}$ is directed upwards, there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ such that $f_{n}\nearrow{\rm ess.sup}_{\mu}\mathcal{A}$ $\mu$-almost surely. By (A.1), we have that ${\rm ess.sup}_{\mu}\mathcal{A}=I_{\min}$ $\mu$-almost surely. Hence, by monotone convergence,
@@
Suppose that $\psi\colon{L}(S)\to(-\infty,\infty]$ satisfies $\psi(0)\in\mathbb{R}$, (T), (M) and is countably maxitive. Then
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\mbox{for all }f\in{L}(S).$
@@
Since $\psi(0)\in\mathbb{R}$, it follows from (T) that the restriction of $\psi$ to $C_{b}(S)$ is real-valued. Hence, it follows from Theorem A.2 that
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\mbox{for all }f\in C_{b}(S).$
Now, let $f\in L(S)$ being bounded from below. Then there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f_{n}\nearrow f$. For instance, $f_{n}(x):=\inf\left\{f(y)+nd(x,y)\colon y\in S\right\}\wedge n$ is as desired. Then since $\psi$ is continuous from below,
@@
Finally, for general $f\in L(S)$, since $\psi$ is countably maxitive, for every $n\in\mathbb{N}$, we obtain
@@
Letting $n\to\infty$, we obtain $\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}$. ∎
@@ Appendix B Auxiliary results for quantile functions
If $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, then there exists a uniformly distributed random variable $U$ with values in $(0,1)$ defined on $(\Omega,\mathcal{F},\mathbb{P})$, see e.g. [3, Theorem 1] and [9, Proposition A.27]. Then $\xi=q(U)$ is as desired. To that end, define $F\colon\mathbb{R}\to[0,1]$ by
@@
see e.g. [21, Theorem 12A]. By replacing $q$ by its left-continuous modification, we can assume that $q$ is left-continuous. Using that $Q(1^{-})=0$, we get that

### v3

@@ 5.5. Dispersive stochastic order
The dispersive order is not antisymmetric. However, by considering the quotient space $\mathcal{M}/_{\sim}$ w.r.t. the equivalence relation $\xi{\sim}\eta$ if and only if $\xi=\eta+c$ for some $c\in\mathbb{R}$, we obtain that the partially ordered space $(\mathcal{M}/_{\sim},\preccurlyeq_{\rm disp})$ is a lattice, see [22].
For a non-empty set $Q$ of quantile functions and $[u,v]\subset[0,1]$, we define its total variation on $[u,v]$ as
${\rm TV}_{[u,v]}\left(Q\right):=\sup_{\pi\in\Pi[u,v]}S_{\pi}(Q),$
where $\Pi[u,v]$ consists of all partitions $\pi=\{u=t_{0}<t_{1}<\cdots<t_{n}=v\}$ for some $n\in\mathbb{N}$. Here, we set $S_{\pi}(Q):=\sum_{i=1}^{n}\sup_{q\in Q}(q(t_{i})-q(t_{i-1}))$ with $q(0):=q(0^{+})$ and $q(1):=q(1^{-})$. We say that the set $Q$ is of bounded variation if ${\rm TV}_{[0,1]}(Q)<\infty$. It is straightforward to verify that ${\rm TV}_{[0,v]}={\rm TV}_{[0,u]}+{\rm TV}_{[u,v]}$ for all $10<u\leq v<$ 1.
Proposition 5.10.
Suppose that $Q:=\{q_{\xi}\colon\xi\in A\}$ is of bounded variation for some $A\subset\mathcal{M}$. Then, there exists $\xi^{\ast}\in\mathcal{M}$ such that $q_{\xi^{\ast}}(t)={\rm TV}_{[0,t]}(Q)$ for all $t\in(0,1]$. Moreover, $\xi^{\ast}$ is a supremum of $A$ w.r.t. $\preccurlyeq_{\rm disp}$.
Proof.
Corollary 5.11.
Proof.
Hence, $A_{0}$ is of bounded variation and has a supremum in $\mathcal{M}_{\infty}$ w.r.t. $\preccurlyeq_{\rm disp}$ by Proposition 5.10. Then passing to the quotient, $A$ has a unique supremum in $\mathcal{M}_{\infty}/_{\sim}$. ∎
It follows from Proposition 5.10 that every sequence $(\xi_{n})_{n\in\mathbb{N}}\subset\mathcal{M}_{\infty}$ which is bounded from above w.r.t. $\preccurlyeq_{\rm disp}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determined by
@@
In the following, we consider the family $\Phi_{\rm disp}:=\{\phi_{u,v}\colon 0<u<v<1\}$ of functions $\phi_{u,v}\colon\mathcal{M}_{\infty}\to\mathbb{R}$ defined by $\phi_{u,v}(\xi):=q_{\xi}(v)-q_{\xi}^{+}(u)$. We identify $\Phi_{\rm disp}$ with the set $\Delta:=\{(u,v)\in\mathbb{R}^{2}\colon 0<u<v<1\}$ which we endow with the Euclidean topology. [^24] Although $\Delta$ is not a complete metric space, it is completely metrizable as it is an open subset, see [33, Theorem 24.12]. In particular, $\Delta$ is a Polish space with the Euclidean topology. It follows from (5.1) that the family $\Phi_{\rm disp}$ induces the dispersive order on $\mathcal{M}_{\infty}$. By Proposition B.1, the mapping $(u,v)\mapsto\phi_{u,v}(\xi)$ is lower semicontinuous on $\Delta$ for all $\xi\in\mathcal{M}_{\infty}$. Hence, the ordered space $(\mathcal{M}_{\infty},\preccurlyeq_{\rm disp})$ satisfies the assumptions of Section 4.
@@
By Proposition 5.10, the sequence $(\xi_{n})_{n\in\mathbb{N}}$ has a supremum $\xi^{\ast}\in\mathcal{M}_{\infty}$ which is determined by
@@
Let $I:=\{s\in\mathbb{R}\colon A_{s}\neq\emptyset\}$. If $\psi$ has countably upwards directed sublevel sets, it follows from Lemma 5.13 that for every $s\in I$, the set $Q_{s}:=\{q_{\xi}\colon\xi\in A_{s}\}$ is of bounded variation, and $\alpha_{\min}(s,u,v)=\beta(s,v)-\beta(s,u)$ for all $(u,v)\in\Delta$ where $\beta(s,u):={\rm TV}_{[0,u]}(Q_{s})$. Moreover, the mapping $\beta\colon\mathbb{R}\times(0,1)\to\mathbb{R}$ is non-decreasing in the second argument and satisfies $\beta(s,v)-\beta(s,u)\leq\beta(s^{\prime},v)-\beta(s^{\prime},u)$ for all $s\leq s^{\prime}$. As a direct application of Theorem 4.1, we obtain the following representation result for maxitive functions w.r.t. the dispersive stochastic order.
Corollary 5.14.
Let $\psi\colon\mathcal{M}_{\infty}\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$. Then, the following conditions are equivalent:
@@
    $\psi$ is monotone and $(A_{s})_{s\in\mathbb{R}}$ are countably upwards directed w.r.t. $\preccurlyeq_{\rm disp}$.
@@
We conclude with an example of a maxitive function for the dispersive order.
Example 5.15.
Consider the function $\beta\colon[0,\infty)\times(0,1)\to\mathbb{R}$, defined by $\beta(s,u):=su$. The function
$\psi(\xi):=\sup_{(u,v)\in\Delta}\frac{q_{\xi}(v)-q_{\xi}(u^{+})}{v-u}\in[0,\infty]$
satisfies the equivalent conditions of Corollary 5.14 and is therefore bounded countably maxitive on $\mathcal{M}_{\infty}/_{\sim}$. The value $\psi(\xi)$ corresponds to the Lipschitz constant of $q_{\xi}$. In particular, if $q_{\xi}$ is differentiable, it holds that
$\psi(\xi)=\sup_{s\in(0,1)}q_{\xi}^{\prime}(s).$
In reliability analysis, the derivative $q_{\xi}^{\prime}(s)$ is referred to as the quantile density function and measures the rate of change of the quantile function; see [30]. When $\xi$ models the uncertain time to failure of a system, $q_{\xi}(u)$ corresponds to the time at which the survival probability of the system is $1-u$. A high value of $q_{\xi}^{\prime}(u)$ implies that small reductions in survival probability correspond to larger increases in time, indicating a slower rate of reliability decline over time. Consequently, $\psi(\xi)$ acts as a measure of the speed at which the reliability of the system decreases, capturing the steepest rate of decline.
6. Conclusion
Maxitive functions are widely used in decision-making under risk and uncertainty, particularly for worst-case or best-case evaluations. In this paper, we extend the concept of maxitive functions to a broad class of preorders, providing a unified framework that goes beyond the classical pointwise order. Specifically, our results address the main stochastic orders in the literature, including the usual stochastic order, the increasing convex order, and the dispersive order.
By transitioning from the pointwise order to weaker stochastic orders, we develop a framework that covers a broader class of maxitive functions. Our main result is a representation theorem for bounded countably maxitive functions. In the translation-invariant case, we obtain representation results in terms of penalized worst-case evaluations. Notably, this generalization also includes widely used decision-making tools such as the value at risk.
@@ Appendix A Auxiliary results for maxitive risk measures on $C_{b}(S)$
The function $\psi$ is continuous from below if $\psi(f)=\lim_{n\to\infty}\psi(f_{n})$ for all sequences $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{X}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ and $f=\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{X}$.
@@
$\psi^{\ast}_{C_{b}(S)}(\mu):=\sup_{f\in C_{b}(S)}\left\{\int f\,{\rm d}\mu-\psi(f)\right\},$
where ${\rm ca}_{1}^{+}(S)$ denotes the set of all Borel probability measures on $S$. In the particular, for the Dirac measure $\delta_{x}$, we get
$I_{\min}(x):=\psi^{\ast}_{C_{b}(S)}(\delta_{x})=\sup_{f\in C_{b}(S)}\{f(x)-\psi(f)\}\quad\mbox{for all }x\in S.$
The following result was recently shown by Delbaen [7, 8]; see also [18] for an alternative proof.
@@
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T), (M) and (C). Then, the following statements are equivalent:
@@
In combination with [15, Theorem 2.1], we obtain the following result. Whereas the representation result in [15] relies on continuity from above, here we rely on the weaker condition of continuity from below, which is guaranteed by countable maxitivity.
@@
Let $\psi\colon C_{b}(S)\to\mathbb{R}$ be a function satisfying (T) and (M). Then, the following statements are equivalent:
@@
    $\psi$ is countably maxitive, i.e., $\psi(f)=\sup_{n\in\mathbb{N}}\psi(f_{n})$ for every $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f=\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$.
@@
    $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$ for every $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$.
@@
The argumentation is similar to that in [15, Theorem 2.1]. Similar to the proof of [15, Theorem 2.1], it holds
@@
for all $\mu\in{\rm ca}_{1}^{+}(S)$, where ${\rm ess.sup}_{\mu}\mathcal{A}$ denotes the essential supremum of $\mathcal{A}$ with respect to the probability measure $\mu$.
$(i)\Rightarrow(ii)$: For $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$, we have
$\psi\Big{(}\sup_{n\in\mathbb{N}}f_{n}\Big{)}=\sup_{n\in\mathbb{N}}\psi(f_{n})\leq 0,$
and therefore $\sup_{n\in\mathbb{N}}f_{n}\in\mathcal{A}$.
$(ii)\Rightarrow(i)$: Let $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ with $f:=\sup_{n\in\mathbb{N}}f_{n}\in C_{b}(S)$. Since by (T), $f_{n}-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$ for all $n\in\mathbb{N}$, we obtain that $f-\sup_{m\in\mathbb{N}}\psi(f_{m})\in\mathcal{A}$. By (T) and (M), it holds
@@
$(i)\Rightarrow(iv):$ Condition (i) ensures that $\psi$ is (finitely) maxitive and continuous from below. In particular, since $\psi$ is maxitive, it satisfies (C) due to [15, Proposition 2.1]. Fix $\mu\in{\rm ca}_{1}^{+}(S)$. Since $\mathcal{A}$ is directed upwards, there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset\mathcal{A}$ with $f_{n}\leq f_{n+1}$ for all $n\in\mathbb{N}$ such that $f_{n}\nearrow{\rm ess.sup}_{\mu}\mathcal{A}$ $\mu$-almost surely. By (A.1), we have that ${\rm ess.sup}_{\mu}\mathcal{A}=I_{\min}$ $\mu$-almost surely. Hence, by monotone convergence,
@@
Suppose that $\psi\colon{L}(S)\to(-\infty,\infty]$ satisfies $\psi(0)\in\mathbb{R}$, properties (T) and (M), and that $\psi(f)=\sup_{n\in\mathbb{N}}\psi(f_{n})$ for every sequence $(f_{n})_{n\in\mathbb{N}}\subset L(S)$ such that $f=\sup_{n\in\mathbb{N}}f_{n}\in L(S)$. Then,
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\text{for all }f\in{L}(S).$
@@
Since $\psi(0)\in\mathbb{R}$, property (T) implies that the restriction of $\psi$ to $C_{b}(S)$ is real-valued. Thus, by Theorem A.2, we have
$\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}\quad\text{for all }f\in C_{b}(S).$
Now, let $f\in L(S)$ be bounded from below. In this case, there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset C_{b}(S)$ such that $f_{n}\nearrow f$. For instance, the sequence defined by
$f_{n}(x):=\inf\left\{f(y)+nd(x,y):y\in S\right\}\wedge n$
satisfies this property. Since $\psi$ is continuous from below, we have
@@
Finally, consider a general $f\in L(S)$. Since $\psi$ is maxitive, for every $n\in\mathbb{N}$, we have
@@
Letting $n\to\infty$, we obtain that $\psi(f)=\sup_{x\in S}\{f(x)-I_{\min}(x)\}$. ∎
@@ Appendix B Auxiliary results for quantile functions
If $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, then there exists a uniformly distributed random variable $U$ with values in $(0,1)$ defined on $(\Omega,\mathcal{F},\mathbb{P})$, see e.g. [6, Theorem 1] and [12, Proposition A.27]. Then $\xi=q(U)$ is as desired. To that end, define $F\colon\mathbb{R}\to[0,1]$ by
@@
see e.g. [26, Theorem 12A]. By replacing $q$ by its left-continuous modification, we can assume that $q$ is left-continuous. Using that $Q(1^{-})=0$, we get that